# Synesthesia of Machines (SoM)-Enhanced ISAC Precoding for Vehicular Networks with Double Dynamics

### v1

@@ Abstract
Integrated sensing and communication (ISAC) technology plays a crucial role in vehicular networks. However, the communication channel within this context exhibits time-varying characteristics, and potential targets may move rapidly, resulting in double dynamics. These presents significant challenges for real-time ISAC precoding design that have not been thoroughly explored. While optimization-based precoding methods have been extensively studied, they are computationally complex and heavily rely on perfect prior information that is rarely available in situations with double dynamics. In this paper, we propose a synesthesia of machine (SoM)-enhanced precoding paradigm, where the base station leverages various modalities such as positioning and channel information to adapt to double dynamics, and effectively utilizes environmental information to stretch ISAC performance boundaries through a deep reinforcement learning framework. Additionally, a parameter-shared actor-critic architecture is tailored to expedite training in complex state and action spaces. Extensive experimental validation has demonstrated the multifaceted superiority of our method over existing approaches.
@@ I Introduction
However, the dynamic nature of the road environment introduces both time-varying characteristics in the wireless channel and rapid movement of the targets, presenting a doubly-dynamic challenge for ISAC. To deal with the time-varying channels, [11] proposed a compressive sensing (CS)-based framework with specialized pilot, reducing the time overhead in channel state information (CSI) acquisition. [12, 13] designed pre-compensators to combat the Doppler effect, mitigating the inter-carrier interference (ICI) in wideband systems. [14] further decoupled the Doppler shifts in angle domain by precoding to effectively suppress the time-variation of communications. In terms of sensing the fast-moving targets, [15] utilized CSI to estimate the target’s velocities, and [16] designed Space-Delay-Doppler precoding by leveraging ICI to enhance the sensing accuracy. Meanwhile, to address the carrier frequency offset resulting from the sensing dynamics, [17] designed a Kalman filter (KF)-based sensing scheme using the estimated Doppler as a prior to filter out time-varying noise and enhance sensing accuracy.
@@
Nevertheless, implementing ISAC under double dynamics remains unexplored and presents unique challenges. Firstly, the complex correlation between the dynamics of communications and sensing can lead to degraded spectrum efficiency (SE) and sensing accuracy. Secondly, existing optimization-based methods are complex and result in high latency, while the need for perfect environmental information such as instantaneous CSI and real-time target positions can be difficult to acquire in double dynamics.
@@
In an effort to lower the dependence on instantaneous state information and mitigate the complexity issue, [18, 19] predicted the real-time environmental prior according to the historical positions, via particle filter (PF) and message passing (MP) respectively. The operating protocol have also been modified for facilitating the predictive precoding. However, these solutions still struggle to address the doubly-dynamic challenge, as they lack robustness in managing the mutual interference between communication and sensing. Moreover, their complexity escalates drastically with a growing number of users, thereby reducing the timeliness of ISAC systems. Additionally, these solutions assume perfect alignment between communication users and sensing targets, which may not be a realistic assumption in real-world doubly-dynamic scenarios. Although obtaining perfect channel-related parameters in double dynamics is challenging, other observable environmental information correlates with channel information, as demonstrated in synesthesia of machine (SoM) [20]. For instance, [21] integrated multi-modal sensing information to boost beamforming for communications. Inspired by this concept, data from diverse modalities has the potential in enhancing the performance of real-time ISAC precoding. However, research from this perspective is still in its early stages, and specific SoM operational modes for ISAC remain undefined.
@@
Notation: $a$, $\bm{a}$ and $\bm{A}$ represent a scalar, a vector and a matrix respectively. $(\cdot)^{\text{T}}$, $(\cdot)^{\text{H}}$, $\text{Tr}(\cdot)$, $\text{rank}(\cdot)$, $\left\|\cdot\right\|_{2}$ and $\left\|\cdot\right\|_{\text{\text{F}}}$ denote transpose, conjugate transpose, trace, rank, 2-norm and Frobenius norm, respectively. $\mathbb{E}(\cdot)$ denotes the expectation. $\mathbb{P}(\cdot)$ denotes the probability. $\mathcal{CN}(m,\sigma^{2})$ represents the complex Gaussian distribution whose mean is $m$ and covariance is $\sigma^{2}$. $\mathbb{R}$ and $\mathbb{C}$ denote the set of real numbers and complex numbers respectively. $|z|$ and $\angle z$ denote the modulus and the phase of a complex number $z$ respectively. $\cap$ and $\cup$ denote the intersection and the union, and $\varnothing$ denotes an empty set.
@@ II-A Transmitting Signal
At the $l$-th symbol duration in the $n$-th subframe ($l=1,\cdots,L$, $n=1,\cdots,T$), the communication symbols for the $U$ users at the $m$-th subcarrier is denoted as $\bm{s}_{\text{m}}^{(n,l)}=\left[s_{\text{m,1}}^{(n,l)},\cdots,s_{\text{m,U}}^{(n,l)}\right]^{\text{T}}$ with $\mathbb{E}\left[\bm{s}_{\text{m}}\bm{s}_{\text{m}}^{\text{H}}\right]=\bm{I}_{\text{U}}$. Specifically, $\bm{s}=\left[\bm{s}_{1}^{\text{T}},\cdots,\bm{s}_{\text{M}}^{\text{T}}\right]^{\text{T}}$ is processed by $M$ frequency-dependent digital precoders $\bm{F}_{\text{BB,m}}\in\mathbb{C}^{N_{\text{RF}}\times U}$ $(m=1,2,\cdots,M)$ respectively, followed by $N_{\text{RF}}$ $M$-point inverse fast Fourier transforms (IFFT). The the $L_{\text{CP}}$-length cyclic prefix (CP) is then added on the time-domain signal, before the time-domain analog precoder $\bm{F}_{\text{RF}}\in\mathbb{C}^{N_{t}\times N_{\text{RF}}}$, which is shared among all the subcarriers. Thus the discrete-time transmitted signal at the $m$-th subcarrier from the array in the $n$-th subframe can be expressed as
$\bm{x}_{\text{m}}^{(n,l)}=\sqrt{\frac{P_{\text{t}}}{M}}\bm{F}_{\text{RF}}^{(n)}\bm{F}_{\text{BB,m}}^{(n)}\bm{s}_{\text{m}}^{(n,l)}=\sqrt{\frac{P_{\text{t}}}{M}}\bm{F}_{\text{m}}^{(n)}\bm{s}_{\text{m}}^{(n,l)},\vspace{-0.2cm}$ (1)
where $P_{\text{t}}$ denotes the total transmitting power. The digital precoder $\bm{F}_{\text{BB,m}}$ is set as $\text{diag}(\left[p_{\text{m,1}},\cdots,p_{\text{m,U}}\right])$, used for power allocation among users. Since the analog precoder $\bm{F}_{\text{RF}}$ is realized by the analog phase shifters (APS), it is constant-modulus. Specifically, when the $B$-bit APS is adopted, we have
$\angle[\bm{F}_{\text{RF}}]_{i,j}\in\mathcal{B}\triangleq\left\{\frac{2\pi b}{2^{B}}|b=0,1,2,\cdots,2^{B}-1\right\}.\vspace{-0.2cm}$ (2)
@@ II-B Communication Model
$\widetilde{\bm{h}}_{\text{u,d}}(l)=\sqrt{\frac{N_{\text{t}}}{P_{\text{u}}}}\sum_{p=1}^{P_{\text{u}}}\beta_{\text{u,p}}g_{\text{rp}}(dT_{\text{s}}-\tau_{\text{u,p}})e^{j\omega_{\text{u,p}}l}\bm{a}_{\text{t}}(\theta_{\text{u,p}}),\vspace{-0.2cm}$ (3)
where $P_{u}$ denotes the number of resolvable paths, $T_{s}$ is the sampling period, $\beta_{u,p}\sim\mathcal{CN}(0,\sigma_{\beta}^{2})$ is the path’s complex gain varying across symbols and $\theta_{u,p}$ is the angle of departure (AoD) of the $p$-th path for the $u$-th user, and $[\bm{a}_{\text{t}}(\theta)]_{\text{i}}=\frac{1}{\sqrt{N_{\text{t}}}}e^{-j(i-1)\pi\sin{\theta}}$ $(i=1,\cdots,N_{\text{t}})$. $g_{\text{rp}(\cdot)}$ is the pulse shaping filter; $\tau_{\text{u,p}}\sim\mathcal{U}(0,(N_{\text{d}}-1)T_{\text{s}})$ is the propagation delay. Denote the carrier frequency as $f_{\text{c}}$, the velocity of light as $c_{\text{v}}$, and the relative velocity of each path as $\bm{v}_{\text{u,p}}$, then the normalized Doppler shift is $\omega_{\text{u,p}}=2\pi f_{\text{c}}|\bm{v}_{\text{u,p}}|T_{\text{s}}\sin{\theta_{\text{u,p}}}/c_{\text{v}}$. Consequently the multi-user channel in time domain is
$\widetilde{\bm{H}}_{\text{d}}(l)=\left[\widetilde{\bm{h}}_{\text{1,d}}^{\text{T}}(l),\widetilde{\bm{h}}_{\text{2,d}}^{\text{T}}(l),\cdots,\widetilde{\bm{h}}_{\text{U,d}}^{\text{T}}(l)\right]^{\text{T}}\in\mathbb{C}^{U\times N_{\text{t}}}.\vspace{-0.2cm}$ (4)
@@
$\bm{H}_{\text{m}}[k]=\frac{1}{M}\sum_{i=0}^{M-1}\sum_{d=0}^{N_{\text{d}}-1}\widetilde{\bm{H}}_{\text{d}}(i+L_{\text{CP}})e^{-j2\pi\frac{kd+(m-k)i}{M}}.\vspace{-0.2cm}$ (5)
@@
$\bm{y}_{\text{c}}=\left[\bm{y}_{\text{1}}^{\text{T}},\bm{y}_{\text{2}}^{\text{T}},\cdots,\bm{y}_{\text{M}}^{\text{T}}\right]^{\text{T}}$ (6)
$=\!\sqrt{\frac{P_{\text{t}}}{M}}\!\underbrace{\left[\begin{matrix}\bm{H}_{\text{1}}[1]\!&\!\cdots\!&\!\bm{H}_{\text{1}}[M]\\ \vdots\!&\!\ddots\!&\!\vdots\\ \bm{H}_{\text{M}}[1]\!&\!\cdots\!&\!\bm{H}_{\text{M}}[M]\end{matrix}\right]}_{\overline{\bm{H}}}\underbrace{\left[\begin{matrix}\bm{F}_{\text{1}}\!&\!\!&\!\\ &\!\ddots\!&\!\\ &\!\!&\!\bm{F}_{\text{M}}\\ \end{matrix}\right]}_{\overline{\bm{F}}}\left[\begin{matrix}\bm{s}_{\text{1}}\\ \vdots\\ \bm{s}_{\text{M}}\end{matrix}\right]\!\!+\!\!\bm{n}_{\text{c}},$
where $\bm{n}_{\text{c}}\sim\mathcal{CN}(\bm{0},\sigma_{\text{c}}^{2}\bm{I}_{\text{UM}})$ is the additive white Gaussian noise (AWGN).
@@
$\widetilde{\bm{h}}_{\text{u,d}}(l)\in\mathcal{C}^{1\times N_{\text{t}}}$ Communication channel per user in time domain
$\widetilde{\bm{H}}_{\text{d}}(l)\in\mathcal{C}^{U\times N_{\text{t}}}$ Overall communication channel in time domain
$\bm{H}_{\text{m}}[k]\in\mathbb{C}^{U\times N_{\text{t}}}$ Inter-carrier response in frequency domain
@@
$\bm{G}_{\text{m}}\in\mathbb{C}^{N_{\text{t}}\times N_{\text{t}}}$ Target response at the $m$-th subcarrier
@@
$\bm{F}_{\text{m}}\in\mathbb{C}^{N_{\text{t}}\times U}$ Equivalent digital precoding matrix
@@
$\bm{h}_{\text{m,k,u}}=\bm{H}_{\text{m}}[k][u,:]$ Frequency-domain inter-carrier channel per user
$\bm{f}_{\text{m,u}}=\bm{F}_{\text{m}}[:,u]$ Equivalent precoding per user per subcarrier
@@
$\!\frac{|\bm{h}_{\text{m,m,u}}^{(n,l)}\bm{f}_{\text{m,u}}^{(n)}|^{2}}{\sum_{k\neq m}\!\sum_{u=1}^{U}\!|\bm{h}_{\text{m,k,u}}^{(n,l)}\bm{f}_{\text{k,u}}^{(n)}\!|^{2}\!+\!\sum_{i\neq u}\!|\bm{h}_{\text{m,m,u}}^{(n,l)}\bm{f}_{\text{m,i}}^{(n)}\!|^{2}\!+\!\frac{N_{\text{t}}UM\sigma_{\text{c}}^{2}}{P_{\text{t}}}},$ (7)
where $\bm{f}_{\text{m,u}}=\bm{F}_{m}[:,u]$ and $\bm{h}_{\text{m,k,u}}=\bm{H}_{\text{m}}[k][u,:]$. Consequently, the spectrum efficiency (SE) can be obtained as

### v2

@@ Abstract
Integrated sensing and communication (ISAC) technology is vital for vehicular networks, yet the time-varying communication channels and rapid movement of targets present significant challenges for real-time precoding design. Traditional optimization-based methods are computationally complex and depend on perfect prior information, which is often unavailable in double-dynamic scenarios. In this paper, we propose a synesthesia of machine (SoM)-enhanced precoding paradigm that leverages modalities such as positioning and channel information to adapt to these dynamics. Utilizing a deep reinforcement learning (DRL) framework, our approach pushes ISAC performance boundaries. We also introduce a parameter-shared actor-critic architecture to accelerate training in complex state and action spaces. Extensive experiments validate the superiority of our method over existing approaches.
@@ I Introduction
However, the dynamic nature of the road environment introduces both time-varying characteristics in the wireless channel and rapid movement of the targets, presenting a doubly-dynamic challenge for ISAC. To deal with the time-varying channels, [11] proposed a compressive sensing (CS)-based framework with specialized pilot, reducing the time overhead in channel state information (CSI) acquisition. [12, 13, 14] designed pre-compensators to combat the Doppler effect, mitigating the inter-carrier interference (ICI) in wideband systems. [15] further decoupled the Doppler shifts in angle domain by precoding to effectively suppress the time-variation of communications. In terms of sensing the fast-moving targets, [16] utilized CSI to estimate the target’s velocities, and [17] designed space-delay-doppler precoding by leveraging ICI to enhance the sensing accuracy. Meanwhile, to address the carrier frequency offset resulting from the sensing dynamics, [18] designed a Kalman filter (KF)-based sensing scheme using the estimated Doppler as a prior to filter out time-varying noise and enhance sensing accuracy.
@@
Nevertheless, implementing ISAC under double dynamics remains unexplored and presents unique challenges. Firstly, the intricate correlation and differences between the dynamics of communication and sensing complicate the simultaneous tracking of both users and targets, resulting in degraded spectrum efficiency (SE) and sensing accuracy. Secondly, existing optimization-based methods are complex and result in high latency. Additionally, the requirement for perfect environmental information, such as instantaneous CSI and real-time target positions, can be challenging to fulfill in dynamic environments.
@@
In an effort to lower the dependence on instantaneous state information and mitigate the complexity issue, [19, 20] predicted the real-time environmental prior according to the historical positions, via particle filter (PF) and message passing (MP) respectively. The operating protocol have also been modified for facilitating the predictive precoding. However, these solutions still struggle to address the doubly-dynamic challenge, as they lack robustness in managing the mutual interference between communication and sensing. Moreover, their complexity escalates drastically with a growing number of users, thereby reducing the timeliness of ISAC systems. Additionally, these solutions assume perfect alignment between communication users and sensing targets, which may not be a realistic assumption in real-world doubly-dynamic scenarios.
Although obtaining perfect channel-related prior in double dynamics is challenging, other observable environmental information correlates with channel information, as demonstrated in synesthesia of machine (SoM) [21]. For instance, [22] predicted the real-time path loss in communication channels based on multi-modal non-CSI data, and [23] integrated multi-modal sensing information to boost beamforming for communications. However, research from this perspective is still in its early stages, and specific SoM operational modes for ISAC remain undefined. The key is to effectively utilize the observations from diverse modalities, such as position information and long-timescale CSI, which can be obtained with significantly lower overhead, to assist in real-time ISAC precoding design. Unlike the location-assisted designs in [24, 25], the inherent sensing capability of the BS can be leveraged to achieve simultaneous communication and positioning, eliminating the need for additional sensors.
@@
Notation: $a$, $\bm{a}$ and $\bm{A}$ represent a scalar, a vector and a matrix respectively. $(\cdot)^{\text{T}}$, $(\cdot)^{\text{H}}$, $\text{Tr}(\cdot)$, $\text{rank}(\cdot)$, $\left\|\cdot\right\|_{2}$ and $\left\|\cdot\right\|_{\text{\text{F}}}$ denote transpose, conjugate transpose, trace, rank, 2-norm and Frobenius norm, respectively. $\mathbb{E}(\cdot)$ denotes the expectation. $\mathbb{P}(\cdot)$ denotes the probability. $\mathcal{CN}(m,\sigma^{2})$ represents the complex Gaussian distribution whose mean is $m$ and covariance is $\sigma^{2}$. $\mathbb{R}$ and $\mathbb{C}$ denote the set of real numbers and complex numbers respectively. $|z|$ and $\angle z$ denote the modulus and the phase of a complex number $z$ respectively. $\bigcap$ and $\bigcup$ denote the intersection and the union, and $\emptyset$ denotes an empty set.
@@ II-A Transmitting Signal
At the $l$-th symbol duration in the $n$-th subframe ($l=1,\cdots,L$, $n=1,\cdots,T$), the communication symbols for the $U$ users at the $m$-th subcarrier is denoted as $\bm{s}_{m}^{(n,l)}=\left[s_{m,1}^{(n,l)},\cdots,s_{m,U}^{(n,l)}\right]^{\text{T}}$ with $\mathbb{E}\left[\bm{s}_{m}\bm{s}_{m}^{\text{H}}\right]=\bm{I}_{\text{U}}$. Specifically, $\bm{s}=\left[\bm{s}_{1}^{\text{T}},\cdots,\bm{s}_{\text{M}}^{\text{T}}\right]^{\text{T}}$ is processed by $M$ frequency-dependent digital precoders $\bm{F}_{\text{BB},m}\in\mathbb{C}^{N_{\text{RF}}\times U}$ $(m=1,2,\cdots,M)$ respectively, followed by $N_{\text{RF}}$ $M$-point inverse fast Fourier transforms (IFFT). The the $L_{\text{CP}}$-length cyclic prefix (CP) is then added on the time-domain signal, before the time-domain analog precoder $\bm{F}_{\text{RF}}\in\mathbb{C}^{N_{t}\times N_{\text{RF}}}$ which is shared among all the subcarriers. Thus the discrete-time transmitted signal at the $m$-th subcarrier from the array in the $n$-th subframe can be expressed as
$\bm{x}_{m}^{(n,l)}=\sqrt{P_{\text{t}}}\bm{F}_{\text{RF}}^{(n)}\bm{F}_{\text{BB},m}^{(n)}\bm{s}_{m}^{(n,l)}=\sqrt{P_{\text{t}}}\bm{F}_{m}^{(n)}\bm{s}_{m}^{(n,l)},\vspace{-0.2cm}$ (1)
where $P_{\text{t}}$ denotes the total transmitting power. The digital precoder $\bm{F}_{\text{BB},m}$ is set as $\text{diag}(\left[p_{m,1},\cdots,p_{m,U}\right])$, used for power allocation among users. Since the analog precoder $\bm{F}_{\text{RF}}$ is realized by the phase shifters (PS), it is constant-modulus. Specifically, when the $B$-bit PS is adopted, we have
$\angle(\bm{F}_{\text{RF}}[i,j])\in\mathcal{B}\triangleq\left\{\frac{2\pi b}{2^{B}}|b=0,1,2,\cdots,2^{B}-1\right\}.\vspace{-0.2cm}$ (2)
@@ II-B Communication Model
$\widetilde{\bm{h}}_{u,d}(l)=\sqrt{\frac{N_{\text{t}}}{P_{\text{u}}}}\sum_{p=1}^{P_{\text{u}}}\beta_{u,p}g_{\text{rp}}(dT_{\text{s}}-\tau_{u,p})e^{j\omega_{u,p}l}\bm{a}_{\text{t}}(\theta_{u,p}),\vspace{-0.2cm}$ (3)
where $P_{\text{u}}$ denotes the number of resolvable paths, $T_{\text{s}}$ is the sampling period, $\beta_{u,p}\sim\mathcal{CN}(0,\sigma_{\beta}^{2})$ is the path’s complex gain varying across symbols and $\theta_{u,p}$ is the angle of departure (AoD) of the $p$-th path for the $u$-th user, and $\bm{a}_{\text{t}}(\theta)[i]=\frac{1}{\sqrt{N_{\text{t}}}}e^{-j(i-1)\pi\sin{\theta}}$ $(i=1,\cdots,N_{\text{t}})$. $g_{\text{rp}(\cdot)}$ is the pulse shaping filter; $\tau_{u,p}\sim\mathcal{U}(0,(N_{\text{d}}-1)T_{\text{s}})$ is the propagation delay. Denote the carrier frequency as $f_{\text{c}}$, the velocity of light as $c_{\text{v}}$, and the relative velocity of each path as $\bm{v}_{u,p}$, then the normalized Doppler shift is $\omega_{u,p}=2\pi f_{\text{c}}|\bm{v}_{u,p}|T_{\text{s}}\sin{\theta_{u,p}}/c_{\text{v}}$. Consequently the multi-user channel in time domain is
$\widetilde{\bm{H}}_{d}(l)=\left[\widetilde{\bm{h}}_{1,d}^{\text{T}}(l),\widetilde{\bm{h}}_{2,d}^{\text{T}}(l),\cdots,\widetilde{\bm{h}}_{U,d}^{\text{T}}(l)\right]^{\text{T}}\in\mathbb{C}^{U\times N_{\text{t}}}.\vspace{-0.2cm}$ (4)
@@
$\bm{H}_{m}[k]=\frac{1}{M}\sum_{i=0}^{M-1}\sum_{d=0}^{N_{\text{d}}-1}\widetilde{\bm{H}}_{d}(i+L_{\text{CP}})e^{-j2\pi\frac{kd+(m-k)i}{M}}.\vspace{-0.2cm}$ (5)
@@
$\bm{y}_{\text{c}}=\left[\bm{y}_{1}^{\text{T}},\bm{y}_{2}^{\text{T}},\cdots,\bm{y}_{M}^{\text{T}}\right]^{\text{T}}$ (6)
$=\!\sqrt{P_{\text{t}}}\!\underbrace{\left[\begin{matrix}\bm{H}_{1}[1]\!&\!\cdots\!&\!\bm{H}_{1}[M]\\ \vdots\!&\!\ddots\!&\!\vdots\\ \bm{H}_{M}[1]\!&\!\cdots\!&\!\bm{H}_{M}[M]\end{matrix}\right]}_{\overline{\bm{H}}}\underbrace{\left[\begin{matrix}\bm{F}_{1}\!&\!\!&\!\\ &\!\ddots\!&\!\\ &\!\!&\!\bm{F}_{M}\\ \end{matrix}\right]}_{\overline{\bm{F}}}\left[\begin{matrix}\bm{s}_{1}\\ \vdots\\ \bm{s}_{M}\end{matrix}\right]\!\!+\!\!\bm{n}_{\text{c}},$
where $\bm{n}_{\text{c}}\sim\mathcal{CN}(\bm{0},\sigma_{\text{c}}^{2}\bm{I}_{UM})$ is the additive white Gaussian noise (AWGN).
@@
$\widetilde{\bm{h}}_{u,d}(l)\in\mathbb{C}^{1\times N_{\text{t}}}$ Communication channel per user in time domain
$\widetilde{\bm{H}}_{d}(l)\in\mathbb{C}^{U\times N_{\text{t}}}$ Overall communication channel in time domain
$\bm{H}_{m}[k]\in\mathbb{C}^{U\times N_{\text{t}}}$ Inter-carrier response in frequency domain
@@
$\bm{G}_{m}\in\mathbb{C}^{N_{\text{t}}\times N_{\text{t}}}$ Target response at the $m$-th subcarrier
@@
$\bm{F}_{m}\in\mathbb{C}^{N_{\text{t}}\times U}$ Equivalent digital precoding matrix
@@
$\bm{h}_{m,k,u}\!=\!\bm{H}_{m}[k][u,:]$ Frequency-domain inter-carrier channel per user
$\bm{f}_{m,u}=\bm{F}_{m}[:,u]$ Equivalent precoding per user per subcarrier
@@
$\!\frac{|\bm{h}_{m,m,u}^{(n,l)}\bm{f}_{m,u}^{(n)}|^{2}}{\sum_{k\neq m}\!\sum_{u=1}^{U}\!|\bm{h}_{m,k,u}^{(n,l)}\bm{f}_{k,u}^{(n)}\!|^{2}\!+\!\sum_{i\neq u}\!|\bm{h}_{m,m,u}^{(n,l)}\bm{f}_{m,i}^{(n)}\!|^{2}\!+\!\frac{N_{\text{t}}UM\sigma_{\text{c}}^{2}}{P_{\text{t}}}},$ (7)
where $\bm{f}_{m,u}=\bm{F}_{m}[:,u]$ and $\bm{h}_{m,k,u}=\bm{H}_{m}[k][u,:]$. Consequently, the spectrum efficiency (SE) can be obtained as